OpenStudy (anonymous):

ITS HOMEWORK NOT AN EXAM ! Simplify Use the Pythagorean Theorem to find the missing side of the right triangle with the given measures. c is the hypotenuse. a = 6, b = 10, c = ?

OpenStudy (anonymous):

a2 + b2 = c2

OpenStudy (anonymous):

11.7 8 136

OpenStudy (hunus):

dw:1367352805884:dw| \[a^{2} + b^{2} = c^{2}\] So c^2 = a^2 + b^2 \[c^{2} = 6^{2} + 10^{2}\] \[c = \sqrt{6^{2} + 10^{2}}\]